Scampini Technologies is expected to generate $150 million in free cash flow next year, and FCF is expected to grow at a constant rate of 3% per year indefinitely. Scampini has no debt or preferred stock, and its WACC is 14%. If Scampini has 55 million shares of stock outstanding, what is the stock's value per share? Do not round intermediate calculations. Round your answer to the nearest cent.

Each share of common stock is worth $ ______ , according to the corporate valuation model.
Answer #1

Value = FCF1 / required rate - growth rate

Value = 150 / 0.14 - 0.03

Value = 150 / 0.11

Value of company = 1,363.636364

Value per share = 1,363.636364/ 55

Value per share = $24.79